If the number of employed (E) and the number of unemployed (U) each increase by 10,000 people, what happens to the unemployment rate? Why?

1
Expert's answer
2021-12-06T16:39:43-0500

"u=\frac{U}{L}=\frac{U}{E+U}"

initially:

U=5000

E=10000

"u=\frac{U}{L}=\frac{U}{E+U}=\frac{5000}{5000+10000}=0.33"

U=15000

E=20000

"u=\frac{U}{L}=\frac{U}{E+U}=\frac{10000}{10000+20000}=0.33"

according to the formula will not change